OpenStudy (anonymous):

Solve this system of equations using substitution. Put your answer in ordered pair form, (x, y). x=4y 3x-2y=20

OpenStudy (anonymous):

substitute x = 4y in the second equation 3(4y) - 2y = 20 12y - 2y = 20 10y = 20 y = 20/10 y=2 substituting y=2 in first equation x = 4y x = 4(2) x = 8 solution (8,2)

OpenStudy (anonymous):

u shld try them now